OpenStudy (anonymous):

Evaluate the definite integral. I will write the equation with teh equation box

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty} dx / (x \sqrt{x^2-1})\]

OpenStudy (anonymous):

is it du/u

OpenStudy (jamesj):

No, not quite

OpenStudy (anonymous):

is it a trig integral? its looks like one

OpenStudy (jamesj):

wait, is the integral \[ \int \frac{x}{\sqrt{x^2-1}} \ dx \]?

OpenStudy (anonymous):

no dx is in the numerator and the numerator is x*sqrt of (x^2 -1)

OpenStudy (anonymous):

i mean teh denominaor is x*sqrt of (....)

OpenStudy (anonymous):

may u pplease show work to see how it gets to infinity

OpenStudy (jamesj):

I mean it behaves like constant*1/sqrt(x-1) which does converge.

OpenStudy (jamesj):

So use the substitution u = sqrt(x^2 - 1) and grind it through. I'll tell you the indefinite integral is arctan( some function of x )

OpenStudy (jamesj):

du = x/sqrt(x^2 - 1) dx Hence 1/(x.sqrt(x^2-1) dx = 1/x^2 . x/sqrt(x^2-1) dx = 1/(u^2 + 1) . du Now integrate that and you have arctan(u) Therefore \[ \int_1^\infty \frac{dx}{x\sqrt{x^2-1}} = \left[ \arctan(\sqrt{x^2-1}) \right]_1^\infty = ... \]

OpenStudy (anonymous):

Thank You! This is a good way for me to start